# AgentArch: A Comprehensive Benchmark to Evaluate Agent Architectures in Enterprise

### v1

@@ Abstract
While individual components of agentic architectures have been studied in isolation, there remains limited empirical understanding of how different design dimensions interact within complex multi-agent systems. This study aims to address these gaps by providing a comprehensive enterprise-specific benchmark evaluating 18 distinct agentic configurations across state-of-the-art large language models. We examine four critical agentic system dimensions: orchestration strategy, agent prompt implementation (ReAct versus function calling), memory architecture, and thinking tool integration. Our benchmark reveals significant model-specific architectural preferences that challenge the prevalent one-size-fits-all paradigm in agentic AI systems. It also reveals significant weaknesses in overall agentic performance on enterprise tasks with the highest scoring models achieving a maximum of only 35.3% success on the more complex task and 70.8% on the simpler task. We hope these findings inform the design of future agentic systems by enabling more empirically backed decisions regarding architectural components and model selection.  [^1] GitHub Repository: https://github.com/ServiceNow/AgentArch
@@ 1 Introduction
Large language models (LLMs) have evolved beyond simple text generation to function as autonomous agents capable of iterative decision-making and complex task completion (Yao et al. 2025). This transformation has opened new possibilities for enterprise automation, where agentic AI systems can interact with existing workflows, databases, and business processes to accomplish multi-step objectives with minimal human intervention.
@@
This paper presents a comprehensive evaluation of 18 distinct agentic architectures across state-of-the-art LLMs, systematically examining the interactions between four key architectural dimensions: orchestration strategy (single-agent vs. multi-agent), agent style (function calling vs. ReAct), memory management (complete vs. summary), and thinking tool integration (Yao et al. 2023). We test each architectural combination on two enterprise use cases of varying complexity, providing practical insights for agentic system designers.
@@
2 Related Work
@@
Description Simple, structured workflow for PTO eligibility verification and request processing. Tests basic multi-step reasoning with clear success criteria. Complex workflow mimicking intelligent customer service systems. Handles simple requests automatically while escalating complex issues. Tests classification and escalation decisions.
@@
Agents 3 agents 9 agents
Agent Types Employee Information Retriever Agent, Balance Eligibility Agent, Leave Approval Agent Request Validation Agent, Duplicate Case Detection Agent, Case Creation Agent, Informational Queries Agent, Transactional Queries Agent, Intent and Sentiment Extraction Agent, Entity Extraction Agent, Document Verification Agent, Email Generation Agent
@@
The evaluation of agentic AI systems has emerged as a critical research area, with existing benchmarks examining various components of agent architectures. However, these efforts have predominantly focused on isolated aspects of agentic systems rather than their integrated performance.
Tool calling is perhaps the most fundamental aspect of all agentic systems. Existing tool calling benchmarks (Liu et al. 2024b; Patil et al. 2025; Wu et al. 2024) have revealed significant LLM weaknesses in key aspects of tool calling including tool selection accuracy, tool input accuracy, valid output formatting, and issues with latency and unnecessary repetition of tool calls. Additional benchmarks focus on more complex scenarios finding that LLM performance on tool calling decreases when combined with complex instruction following (Yao et al. 2025) and multi step reasoning in nested tool calling scenarios (Basu et al. 2025).
Another key aspect of agentic systems is memory and context management. LLMs are inherently limited by their context windows which provides a unique challenge for those building agentic systems to ensure that the LLMs keep the most relevant information in context. This is relevant for both conversational and tool-use settings, as conversational context accumulates over time, while tool calls can significantly increase context length due to their frequency or output size (Maharana et al. 2024a). Existing research presents a variety of mechanisms for improving memory management in agentic systems such as treating agent memory as a self directed operating system with different storage tiers (Packer et al. 2024) , creating dynamic knowledge networks for connecting information (Xu et al. 2025), and implementing a dual-track memory system meant to resemble long-term and short-term memory in the human brain (Liu et al. 2024a).
As LLMs enable more complex agentic workflows, Multi-Agent Systems (MAS) have emerged to support use cases involving specialized agents, prompting the development of benchmarks to evaluate their effectiveness. (Zhu et al. 2025) examine LLM performance across various MAS structures, including hierarchical and distributed planning, evaluating task completion and planning capabilities to assess LLM effectiveness in both planning and execution roles. (Agashe, Fan, and Wang 2025) and (Geng and Chang 2025) focus specifically on evaluating planning and coordination between agents in MAS finding that planning abilities are a key component of success in MAS. (Yue et al. 2025) illustrate how multi-agent systems can benefit from routing frameworks that help ensure that the appropriate agents, communication strategies, and LLMs are selected for a given task. Similarly, (Hou, Tang, and Wang 2025) introduce a dynamic MAS framework that incorporates hierarchical planning along with a search process for selecting the optimal multi-agent configuration for the task at hand. Other work on dynamically generated MAS demonstrates how scaling single agents into MAS can lead to performance increases (Yuan et al. 2025). Recent work compares heterogeneous LLMs versus a single model in MAS, highlighting how MAS can be used to leverage individual model strengths (Ye et al. 2025).
As our work is primarily focused on enterprise benchmarks, we also explore existing work on agents in enterprise domains. Enterprise-focused evaluation has consistently revealed a gap between existing academic benchmarks and real-world applicability, with even state-of-the-art models struggling on simple workplace tasks (Styles et al. 2024). Benchmarks across different enterprise domains consistently show that realistic business task complexity significantly exceeds what current models can handle reliably, with performance typically degrading in multi-turn interactions and interconnected tasks (Huang et al. 2025a, b; Drouin et al. 2024; Styles et al. 2024).
While prior benchmarks have advanced understanding of individual agentic components, they remain fragmented—evaluating tool use (Yao et al. 2025; Patil et al. 2025) (Basu et al. 2025), memory (Maharana et al. 2024b), and coordination (Geng and Chang 2025; Agashe, Fan, and Wang 2025) in isolation. Existing enterprise benchmarks address task complexity but lack systematic analysis of how architectural choices interact across scenarios. As a result, developers lack empirical guidance for selecting and combining agentic components under real-world enterprise constraints.
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 36.9 21.2 6.2 0.0 7.1 0.0 1.2 0.0 0.4 0.0 28.1 0.0
Complete Disabled 31.9 17.5 4.8 1.2 7.3 0.5 1.2 0.0 0.0 0.0 28.7 0.2
Function Calling Complete Enabled 58.8 14.8 50.2 4.2 65.8 4.8 4.6 20.5 0.0 0.0 67.7 15.2
Complete Disabled 54.4 16.2 46.5 5.0 56.0 4.8 5.4 21.8 0.0 0.0 57.7 17.2
Summarized Enabled 51.9 15.2 49.0 0.8 59.0 2.5 5.0 18.5 0.0 0.0 59.6 16.8
Summarized Disabled 52.7 15.8 45.0 0.8 49.6 1.0 4.8 19.0 0.0 0.0 44.2 16.5
Orch-Isolated React Complete Enabled 28.7 17.2 3.8 0.0 4.6 0.0 2.3 0.0 0.2 0.0 28.3 0.0
Complete Disabled 25.8 16.0 5.8 0.2 5.6 0.0 1.5 0.0 0.0 0.0 28.1 0.5
Function Calling Complete Enabled 58.8 17.2 53.5 1.8 67.1 0.8 6.0 22.2 0.0 0.0 68.5 14.8
Complete Disabled 54.6 16.2 52.1 1.8 57.1 1.5 4.0 21.0 0.0 0.0 57.5 16.2
Summarized Enabled 54.6 16.5 46.0 1.8 59.6 2.0 2.9 20.5 0.0 0.0 58.1 16.8
Summarized Disabled 51.5 15.8 44.2 3.0 54.4 1.5 4.8 21.2 0.0 0.0 44.8 14.5
Single Agent React Complete Enabled 34.4 1.0 15.8 0.2 51.2 1.2 8.8 1.5 6.2 0.0 63.7 3.5
Complete Disabled 36.7 3.8 9.6 0.0 41.5 1.2 2.9 2.2 12.3 0.0 52.7 9.0
Function Calling Complete Enabled 68.1 22.2 39.0 3.0 35.4 0.2 56.7 1.5 0.0 0.0 66.7 33.8
Complete Disabled 49.6 21.2 26.9 3.5 19.8 0.2 55.8 2.0 3.8 0.0 31.7 35.0
Summarized Enabled 70.8 21.5 37.9 2.5 33.5 0.5 55.0 1.5 0.6 0.0 64.4 33.2
Summarized Disabled 48.5 19.8 28.7 2.8 24.4 0.5 55.8 1.0 0.0 0.0 31.7 35.2
Table 2: Pass@1 Acceptable Rate. Bolded scores indicate highest score per model on use case.
@@ 3.1 Experiment Set Up
Use Cases We evaluate two enterprise workflows of varying complexity to examine how task complexity affects agentic performance on real-world enterprise automation scenarios. Both of these use cases involve specific directions to complete tasks in a specified order, reflecting the rigid nature of many enterprise workflows where steps must take place in a predefined order. For each request in a use case, we define an expected final decision (e.g. leave approve or rejected). We create mock data so that each tool request returns a deterministic result. Unlike existing benchmarks that use simplified, clean responses, we deliberately construct enterprise-realistic data that is complex, lengthy, and messy. Knowledge base articles span thousands of words with extensive documentation reflecting real enterprise systems. Tool responses return complex JSON objects where relevant information is often buried within metadata, error codes, and auxiliary fields. This approach better represents the information processing challenges agents face in production environments where data is rarely clean or perfectly structured.
Figure 1: Orchestration, Isolated Agents. Agents ask orchestrator for help and orchestrator selects agent to assist.
@@
Customer Request Routing (CR) represents a complex workflow that mimics intelligent customer service systems, handling simple requests automatically while escalating complex issues to human agents. This tests nuanced classification, context understanding, and appropriate escalation decisions with 31 custom tools and 9 agents. Each use case includes 60 user utterances designed to cover the full spectrum of real-world requests, including successful scenarios, edge cases, and failure conditions that would occur in actual enterprise deployments. Implementation details are provided in Appendix A.1.
Orchestration Strategies. We evaluate four orchestration strategies to examine single-agent versus multi-agent performance across key dimensions: decision-making with varying option sets, collaboration costs and benefits, and specialist versus generalist agent effectiveness. 1) Orchestrator-led, Isolated Agent. An orchestrator assigns tasks to specialist agents and controls all inter-agent communication. When agents need additional information, they request help through the orchestrator, who selects an appropriate agent to respond. This tests whether centralized planning and controlled information flow improve outcomes. 2) Orchestrator-led, Open Agent Network. The orchestrator assigns initial tasks, but agents communicate directly with each other when additional information is needed. Agents select collaborators independently using agent descriptions. This examines whether direct agent-to-agent communication is more efficient than orchestrator-mediated interaction. 3) Single Agent. A single agent receives the user request and has access to all available tools, with no collaboration mechanisms. This tests agent performance in large, unfiltered option spaces.
Agents. We evaluate two prevalent agent styles to examine how different prompting paradigms influence model performance. With function calling, the model directly selects tools from the available toolkit, leveraging inherent function calling capabilities without explicit intermediate reasoning. Using ReAct, The model is instructed to output its thoughts before selecting actions thus using a structured reasoning-action framework (Yao et al. 2023). We investigate whether the explicit reasoning in ReAct prompting yields substantially different outcomes compared to direct function calling. Implementation details and prompts are provided in the appendix.
Memory Management. We evaluate two information sharing approaches to examine the trade-off between context completeness and input context length. With complete memory, agents see all previous tool calls, parameters, and responses from other agents, maximizing context but increasing prompt length. With summarized memory, Agents receive only final summaries from previous agents, reducing context length while potentially limiting coordination effectiveness. This tests whether comprehensive information sharing improves performance or whether condensed summaries provide sufficient context for effective collaboration.
Thinking Tools. We evaluate whether providing explicit reasoning tools improves model performance on multi-step analysis tasks. When thinking tools are enabled, models receive access to math for calculations and synthesize_collected_information for data organization, particularly relevant for PTO eligibility where models must calculate date ranges and compare leave balances. We hypothesize that models without built-in reasoning capabilities will benefit from these structured tools. See appendix A.2 for more details and examples.
@@
Model Settings. We evaluate our benchmark using six models: OpenAI GPT-4.1, OpenAI-GPT-4o, LLaMA 3.3 70B, OpenAI GPT-4.1-mini, Anthropic Claude Sonnet 4, and OpenAI o3-mini. We evaluated Claude Sonnet 4 with reasoning disabled. For all experiments, we set temperature to 0 and use no other sampling parameters. We selected this set of models to represent state-of-the-art models of varying sizes, including models with and without in-built reasoning capabilities. For function calling architectures, we use the in-built OpenAI function calling APIs for OpenAI models and LLaMa 3.3. For Sonnet 4, we use the Bedrock API.
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 1.0 20.8 19.4 77.2 15.2 64.8 64.0 96.5 29.2 35.6 0.0 31.0
Complete Disabled 2.1 18.8 22.5 71.5 18.3 65.2 67.5 98.0 28.3 30.1 0.0 26.8
Function Calling Complete Enabled 0.0 0.0 42.9 16.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Complete Disabled 0.0 0.0 58.3 18.8 0.0 0.0 0.0 0.0 0.2 0.0 0.0 0.0
Summarized Enabled 0.0 0.0 49.6 7.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Disabled 0.0 0.0 48.1 4.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Orch-Isolated React Complete Enabled 2.3 13.8 35.0 77.8 11.0 63.2 68.1 96.8 24.4 22.1 0.0 36.2
Complete Disabled 4.4 19.8 31.0 86.8 16.5 62.5 66.2 97.5 23.3 28.7 0.2 31.8
Function Calling Complete Enabled 0.0 0.0 0.4 19.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Complete Disabled 0.0 0.0 0.2 10.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Enabled 0.0 0.0 0.0 8.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Disabled 0.0 0.0 0.0 4.2 0.0 0.0 0.0 0.0 0.2 0.0 0.0 0.0
Single Agent React Complete Enabled 0.0 0.0 0.0 11.8 0.0 0.5 0.0 0.0 0.2 0.0 0.0 0.0
Complete Disabled 1.9 0.2 4.2 8.5 0.2 0.0 0.0 0.0 3.3 0.0 0.0 0.0
Function Calling Complete Enabled 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Complete Disabled 0.0 0.0 0.0 0.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Enabled 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Disabled 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Table 4: Pass@1 Hallucination Rate ↓ Bolded scores indicate highest score (most hallucinations) per model on use case

### v2

# AgentArch: A Benchmark for Evaluating Agent Architectures in Enterprise Workflows

@@ Abstract
While individual components of agentic architectures have been studied in isolation, we lack empirical evidence on how design dimensions interact in realistic multi-agent systems. This study aims to address these gaps by providing a comprehensive enterprise-specific benchmark. We introduce AgentArch—a benchmark that evaluates 18 agentic configurations across state-of-the-art large language models. The study spans four dimensions: orchestration strategy, agent style (ReAct versus function calling), memory architecture, and thinking tool integration. Results reveal strong model-specific architectural preferences that challenge one-size-fits-all assumptions in agentic AI. They also expose notable performance gaps on enterprise tasks: the best models reach only 35.3% success on the complex task and 70.8% on the simpler task. We aim to inform architecting choices for enterprise agents by grounding model selection and component design in systematic evidence.
AgentArch: A Benchmark for Evaluating Agent Architectures in Enterprise Workflows
Tara Bogavelli, Hari Subramani, Roshnee Sharma ServiceNow Correspondence: tara.bogavelli@servicenow.com https://github.com/ServiceNow/AgentArch
@@ 1 Introduction
Large language models (LLMs) have evolved beyond simple text generation to function as autonomous agents capable of iterative decision-making and complex task completion taubench:25. This transformation has opened new possibilities for enterprise automation, where agentic AI systems can interact with existing workflows, databases, and business processes to accomplish multi-step objectives with minimal human intervention.
@@
This paper presents a comprehensive evaluation of 18 distinct agentic architectures across state-of-the-art LLMs, systematically examining the interactions between four key architectural dimensions: orchestration strategy (single agent vs. multi-agent), agent style (function calling vs. ReAct), memory management (complete vs. summary), and thinking tool integration react:23. We test each architectural combination on two enterprise use cases of varying complexity, providing practical insights for agentic system designers.
@@
To summarize, our work makes the following contributions:
  • We introduce AgentArch, the first benchmark systematically evaluating 18 agentic architectures across six large language models on realistic enterprise workflows.
  • We jointly analyze the effects of orchestration, prompting style, memory design, and thinking tool integration, providing quantitative insight into how these factors interact in practice.
  • Our experiments reveal model-specific architectural preferences, showing that optimal configurations vary by model and task complexity, rather than following a single best-performing design.
  • We offer evidence-based recommendations for selecting models and architectural configurations suited to enterprise reliability and integration constraints.
Overall, AgentArch contributes a systematic, empirically grounded step toward understanding how architectural design choices influence agentic system behavior, supporting more informed and reproducible development practices.
@@
Description Simple workflow for PTO eligibility and request processing. Tests basic multi-step reasoning and counting. Complex workflow mimicking customer service systems. Automates simple requests, escalating complex issues. Tests classification and instruction following.
@@
Agents Employee Information Retriever, Balance Eligibility, Leave Approval Request Validation, Duplicate Case Detection, Case Creation, Informational Queries, Transactional Queries, Intent and Sentiment Extraction, Entity Extraction, Document Verification, Email Generation
@@
2 Related Work
Evaluating agentic AI systems has become a central research focus, yet many studies still assess components in isolation rather than integrated end-to-end performance.
Tool calling. Tool use is foundational for agentic systems. Benchmarks such as agentbench:24; bfcl:25; sealtools:24 document persistent weaknesses in tool selection, argument fidelity, output formatting, latency, and redundant calls. More complex settings further degrade performance: combining tool use with intricate instruction following taubench:25 or multi-step, nested tool invocations basu2025nestfulbenchmarkevaluatingllms increases error rates.
Memory and context management. Limited context windows and growing conversational and tool-response histories create pressure on memory design. Prior work explores tiered memory as an OS abstraction packer2024memgptllmsoperatingsystems, dynamic knowledge graphs for retrieval and linking xu2025amemagenticmemoryllm, and dual-track (short/long-term) schemes liu2024llmconversationalagentmemory. Empirical studies highlight how tool calls inflate context via frequent, verbose outputs maharana2024evaluatinglongtermconversationalmemory, motivating mechanisms that preserve relevance while controlling prompt length.
Multi-Agent Systems (MAS). With increasing workflow complexity, MAS introduce specialization and coordination. Benchmarks examine planning structures (hierarchical vs. distributed) and execution quality marble:25, as well as inter-agent planning/coordination llmcoord:25; geng2025realmbenchrealworldplanningbenchmark, identifying planning reliability as a bottleneck. Routing frameworks improve agent/model selection and communication patterns masrouter:25; adaptive approaches search multi-agent configurations online hou2025halohierarchicalautonomouslogicoriented. Automatic scaling from single agents to MAS can yield gains yuan2025evoagentautomaticmultiagentgeneration, and heterogeneous-model MAS leverage complementary strengths ye2025xmasbuildingmultiagentsystems.
Enterprise evaluations. Enterprise-focused studies consistently reveal a gap between academic benchmarks and real deployments. Even state-of-the-art models struggle on routine workplace tasks workbench:24, with performance degrading in multi-turn and interdependent workflows crmarena:25; huang2025crmarenaproholisticassessmentllm; workarena:24; workbench:24. These findings underscore the need for evaluations that reflect realistic business constraints: reliability, strict procedure orderings, and integration with existing systems.
Prior work has advanced understanding of individual components—tool use taubench:25; bfcl:25; basu2025nestfulbenchmarkevaluatingllms, memory locomo:24, and coordination geng2025realmbenchrealworldplanningbenchmark; llmcoord:25—but remains fragmented. Enterprise-oriented benchmarks capture task complexity yet rarely analyze interactions among orchestration, prompting style, memory strategy, and reasoning aids. Our benchmark fills this gap by jointly varying these dimensions across two enterprise use cases, providing evidence on model-specific architectural preferences and the relative advantages of multi-agent versus single agent designs.
@@ 3.1 Experiment Set Up
Use Cases We evaluate two enterprise workflows of differing complexity to examine how task complexity affects agentic performance on real-world enterprise automation scenarios. Both of these use cases involve highly-specific directions to complete tasks in a given order, reflecting the rigid nature of many enterprise workflows where steps must take place in a predefined order. For each request in a use case, we define an expected final decision (e.g. leave approved or rejected). We create mock data so that each tool request returns a deterministic result. Unlike existing benchmarks that use simplified, clean responses, we deliberately construct enterprise-realistic data that is complex, lengthy, and messy. Knowledge base articles span thousands of words with extensive documentation reflecting real enterprise systems. Tool responses return complex JSON objects where relevant information is often buried within metadata, error codes, and auxiliary fields. This approach better represents the information processing challenges agents face in production environments where data is rarely clean or perfectly structured.
Customer Request Routing (CR) represents a complex workflow that mimics intelligent customer service systems, handling simple requests automatically while escalating complex issues to human agents. This tests nuanced classification, context understanding, and appropriate escalation decisions with 31 custom tools and 9 agents. Each use case includes 60 user utterances designed to cover the full spectrum of real-world requests, including successful scenarios, edge cases, and failure conditions that would occur in actual enterprise deployments. Implementation details are provided in Appendix A.
Orchestration Strategies. We evaluate three orchestration strategies to examine single agent versus multi-agent performance across key dimensions: decision-making with varying option sets, collaboration costs and benefits, and specialist versus generalist agent effectiveness. 1) Orchestrator-led, Isolated Agent. An orchestrator assigns tasks to specialist agents and controls all inter-agent communication. When agents need additional information, they request help through the orchestrator, who selects an appropriate agent to respond. This tests whether centralized planning and controlled information flow improve outcomes. 2) Orchestrator-led, Open Agent Network. The orchestrator assigns initial tasks, but agents communicate directly with each other when additional information is needed. Agents select collaborators independently using agent descriptions. This examines whether direct agent-to-agent communication is more efficient than orchestrator-mediated interaction. 3) Single Agent. A single agent receives the user request and has access to all available tools, with no collaboration mechanisms. This tests agent performance in larger, unfiltered option spaces.
Figure 1: Orchestration, Isolated Agents. Agents ask the orchestrator for help, and the orchestrator selects agent to assist.
Agents. We evaluate two prevalent agent styles to examine how different prompting paradigms influence model performance. With function calling, the model directly selects tools from the available toolkit, leveraging inherent function calling capabilities without explicit intermediate reasoning. Using ReAct, the model outputs its reasoning before selecting actions, following a structured reasoning-action framework react:23. We investigate whether the explicit reasoning in ReAct prompting yields substantially different outcomes compared to direct function calling. Implementation details and prompts are provided in the appendix.
Memory Management. We evaluate two information-sharing approaches to examine the trade-off between context completeness and input context length. With complete memory, agents see all previous tool calls, parameters, and responses from other agents, maximizing context but increasing prompt length. With summarized memory, agents receive only final summaries from previous agents thus reducing context length.This tests whether comprehensive information sharing enhances performance or condensed summaries suffice for effective collaboration.
Thinking Tools. We evaluate whether providing explicit reasoning tools improves model performance on multi-step analysis tasks. When thinking tools are enabled, models receive access to math for calculations and synthesize_collected_information for data organization, particularly relevant for PTO eligibility, where models must calculate date ranges and compare leave balances. We hypothesize that models without built-in reasoning capabilities will benefit from these structured tools. See Appendix A.2 for more details and examples.
@@
Model Settings. We evaluate our benchmark using six models: OpenAI GPT-4.1, OpenAI-GPT-4o, LLaMA 3.3 70B, OpenAI GPT-4.1-mini, Anthropic Claude Sonnet 4, and OpenAI o3-mini. We evaluated Claude Sonnet 4 with reasoning disabled. For all experiments, we set the temperature to 0 and use no other sampling parameters. We selected this set of models to represent state-of-the-art models of varying sizes, including models with and without in-built reasoning capabilities. For function calling architectures, we use the in-built OpenAI function calling APIs for all models besides Sonnet 4 which uses the Bedrock API.
Figure 3: Pass@1 Acceptable Rate. Heatmap cells report the Pass@1 Acceptable score (higher/darker is better) for each model–configuration pair on the two enterprise tasks: Time Off (TO) and Customer Routing (CR). Pass@1 is computed over $k=8$ attempts per configuration. Overall, function-calling configurations tend to outperform ReAct, thinking tools often help on the simpler TO task, and no single architecture dominates across models or tasks.